A ball is gently dropped from a height of 20 m. If its velocity increases uniformly at the rate of 10 m s-2, with what velocity will it strike the ground? After what time will it strike the ground?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Assume, the final velocity with which the ball will strike the ground be ‘v’and time it takes to strike the ground be ‘t&Rsquo;
Initial Velocity of ball, u =0
Distance or height of fall, s =20 m
Downward acceleration, a =10 m s^-2
As we know, 2as =v^2-u^2
v^2= 2as+ u^2
= 2 x 10 x 20 + 0
= 400
∴Final velocity of ball, v = 20 ms^-1
t = (v-u)/a
∴Time taken by the ball to strike = (20-0)/10
= 20/10
= 2 seconds

Thanks